### v2

@@ 4.2 Experimental Setup
On the other hand, the research’s dataset was randomly partitioned into 70% training, 15% validation, and 15% testing sets. In contrast, for a fair comparison, the CatchPhish dataset was divided into 80% training and 20% testing, following Aljofey’s approach [aljofey2022effective].
@@ 4.3 Model Performance
In this section, we present a comprehensive analysis of the performance of our proposed MultiText-LP model across various dimensions. We conduct a meticulous examination by comparing our model against individual pretrained NLP models, a standalone MLP model, and an existing approach, Aljofey’s [aljofey2022effective], on the ’CatchPhish’ dataset. Our goal is to showcase the enhanced classification capabilities achieved through the fusion of NLP and MLP models.
@@
To contextualize the performance of MultiText-LP, we begin by summarizing the performance of various pretrained models on specific features (Table 5). For ’NLP-1 (Page Title),’ the CANINE model outperforms others, boasting an F1-Score of 93.33% and an accuracy of 93.84%. In the ’NLP-2 (Page Content)’ category, RoBERTa demonstrates superior performance with an F1-Score of 96.38% and an accuracy of 96.76%.
@@
Table 7: Approach Performance Comparison on the CatchPhish Dataset
@@
Aljofey’s Approach [aljofey2022effective] 89.00 88.20 87.68 87.94
@@
A detailed evaluation in Table 7 focuses on HTML content within the challenging CatchPhish dataset. Aljofey’s Approach [aljofey2022effective] demonstrates commendable metrics, yet the MultiText-LP model showcases superior results with an accuracy of 89.58%, precision of 89.61%, and F1-Score of 88.17%. Despite a slight adjustment in recall (87.03%), our analysis reveals positive increments in accuracy (+0.58%), precision (+1.41%), and F1-Score (+0.23%). This emphasizes the enhanced efficacy of the MultiText-LP model in HTML content classification compared to Aljofey’s Approach.
Our innovative MultiText-LP model emerges as a powerful tool for HTML content classification, exhibiting superior performance compared to individual NLP models, a standalone MLP model, and a state-of-the-art approach. The fusion of two NLP and MLP demonstrates a synergistic effect, enhancing classification capabilities. With promising results on the challenging CatchPhish dataset, our approach holds significant potential for applications in the broader HTML content analysis and classification field.
@@ 5 Limitations
Another limitation is the limited availability of datasets, which includes HTML content. We can only access the CatchPhish dataset from Aljofey’s study [aljofey2022effective]. This scarcity makes it challenging to find comparable datasets and hinders direct comparisons with other approaches. Moreover, there is a lack of prior work focusing explicitly on dealing with HTML content. This absence of existing methodologies for working solely with HTML further complicates our efforts to benchmark and evaluate our approach compared to established methods.
@@ 6 Conclusion and Future Work
While research efforts have focused on detecting phishing websites, a notable gap exists in analyzing HTML content. Several studies have explored URLs and HTML, but a notable gap exists in research focusing on analyzing HTML content alone. Because of that, there is very limited dataset accessibility. To address this gap, we curated an up-to-date dataset that reflects real-life phishing conditions, ensuring its relevance and applicability. This dataset, shared openly with the community, contributes to the existing literature and provides a valuable resource for comprehensive phishing research.
@@
Our innovative MultiText-LP model emerges as a powerful tool for HTML content classification, exhibiting superior performance compared to individual NLP models, a standalone MLP model, and a state-of-the-art approach. The fusion of two NLP and MLP demonstrates a synergistic effect, enhancing classification capabilities. With promising results on the challenging CatchPhish dataset, our approach holds significant potential for applications in the broader HTML content analysis and classification field.

### v3

@@ 4.2 Experimental Setup
On the other hand, the research’s dataset was randomly partitioned into 70% training, 15% validation, and 15% testing sets. In contrast, for a fair comparison, the Aljofey’s dataset was divided into 80% training and 20% testing, following Aljofey’s approach [16].
@@ 4.3 Model Performance
In this section, we present a comprehensive analysis of the performance of our proposed MultiText-LP model across various dimensions. We conduct a meticulous examination by comparing our model against individual pretrained NLP models, a standalone MLP model, and an existing approach, Aljofey’s [16], on the research dataset. Our goal is to showcase the enhanced classification capabilities achieved through the fusion of NLP and MLP models.
@@
To contextualize the performance of MultiText-LP, we begin by summarizing the performance of various pretrained models on specific features (Table 5). For ‘NLP-1 (Page Title),’ the CANINE model outperforms others, boasting an F1-Score of 93.33% and an accuracy of 93.84%. In the ‘NLP-2 (Page Content)’ category, RoBERTa demonstrates superior performance with an F1-Score of 96.38% and an accuracy of 96.76%.
@@
Table 7: Approach Performance Comparison on the Aljofey’s Dataset
@@
Aljofey’s Approach [16] 89.00 88.20 87.68 87.94
@@
A detailed evaluation in Table 7 focuses on HTML content within the challenging Aljofey’s dataset. Aljofey’s Approach [16] demonstrates commendable metrics, yet the MultiText-LP model showcases superior results with an accuracy of 89.58%, precision of 89.61%, and F1-Score of 88.17%. Despite a slight adjustment in recall (87.03%), our analysis reveals positive increments in accuracy (+0.58%), precision (+1.41%), and F1-Score (+0.23%). This emphasizes the enhanced efficacy of the MultiText-LP model in HTML content classification compared to Aljofey’s Approach.
Our innovative MultiText-LP model emerges as a powerful tool for HTML content classification, exhibiting superior performance compared to individual NLP models, a standalone MLP model, and a state-of-the-art approach. The fusion of two NLP and MLP demonstrates a synergistic effect, enhancing classification capabilities. With promising results on the challenging Aljofey’s dataset, our approach holds significant potential for applications in the broader HTML content analysis and classification field.
@@ 5 Limitations
Another limitation is the limited availability of datasets, which includes HTML content. We can only access the Aljofey’s dataset from Aljofey’s study [16]. This scarcity makes it challenging to find comparable datasets and hinders direct comparisons with other approaches. Moreover, there is a lack of prior work focusing explicitly on dealing with HTML content. This absence of existing methodologies for working solely with HTML further complicates our efforts to benchmark and evaluate our approach compared to established methods.
@@ 6 Conclusion and Future Work
While research efforts have focused on detecting phishing websites, a notable gap exists in analyzing HTML content. Several studies have explored URLs and HTML, but a noteworthy gap exists in research focusing on analyzing HTML content alone. Because of that, dataset accessibility is minimal. To address this gap, we curated an up-to-date dataset that reflects real-life phishing conditions, ensuring its relevance and applicability. This dataset, shared openly with the community, contributes to the existing literature and provides a valuable resource for comprehensive phishing research.
@@
Our innovative MultiText-LP model emerges as a powerful tool for HTML content classification, exhibiting superior performance compared to individual NLP models, a standalone MLP model, and a state-of-the-art approach. The fusion of two NLP and MLP demonstrates a synergistic effect, enhancing classification capabilities. With promising results on the challenging Aljofey’s dataset, our approach holds significant potential for applications in the broader HTML content analysis and classification field.